how much water should be added to 77 gallons of a 32% salt solution to get a 30% salt solution

Answer by robertb(5830)

0.00x + 0.32*77 = 0.30(x + 77)
==> 24.64 = 0.30x + 23.1
==> 1.54 = 0.30x
==> x = 77/15 gallons of water.